OpenStudy (anonymous):

Someone please help log(x^3y^5)-2log(x√3:(y))-3log(x/y)

OpenStudy (anonymous):

Properties of Logarithms \[\log(abc) = \log a + \log b + \log c\] and \[\log \frac{ ab }{ c} = loga + \log b - \log c\] and \[loga ^{x} = xloga\]